# 3D Diffusion Policy

### v1

@@ Abstract
Imitation learning provides an efficient way to teach robots dexterous skills; however, learning complex skills robustly and generalizablely usually consumes large amounts of human demonstrations. To tackle this challenging problem, we present 3D Diffusion Policy (DP3), a novel visual imitation learning approach that incorporates the power of 3D visual representations into diffusion policies, a class of conditional action generative models. The core design of DP3 is the utilization of a compact 3D visual representation, extracted from sparse point clouds with an efficient point encoder. In our experiments involving 72 simulation tasks, DP3 successfully handles most tasks with just 10 demonstrations and surpasses baselines with a 55.3% relative improvement. In 4 real robot tasks, DP3 demonstrates precise control with a high success rate of 85%, given only 40 demonstrations of each task, and shows excellent generalization abilities in diverse aspects, including space, viewpoint, appearance, and instance. Interestingly, in real robot experiments, DP3 rarely violates safety requirements, in contrast to baseline methods which frequently do, necessitating human intervention. Our extensive evaluation highlights the critical importance of 3D representations in real-world robot learning. Videos, code, and data are available on 3d-diffusion-policy.github.io.
@@ I Introduction
Safety. DP3 consistently adheres to all safety requirements in real-world tasks, unlike baseline methods which often give erratic commands and exhibit unexpected behaviors, posing potential damage to the robot hardware.
@@ II-A Diffusion Models in Robotics
Diffusion models, a category of generative models that progressively transform random noise into a data sample, have achieved great success in high-fidelity image generation [24, 64, 52, 63]. Owing to their impressive expressiveness, diffusion models have recently been applied in robotics, including in fields such as reinforcement learning [69, 2], imitation learning [10, 41, 51, 71], reward learning [26, 39], grasping [70, 66, 62], and motion planning [53, 28]. In this work, we focus on representing visuomotor policies as conditional diffusion models, referred to as diffusion policies, following the framework established in [10, 41]. Unlike prior methods that primarily focus on images and states as conditions, we pioneer in incorporating 3D conditioning into diffusion policies.
@@ III-B Perception
| Algorithm Task | Hammer | Door | Pen | Assembly | Basketball | Shelf Place | Laptop | Faucet | Toilet | Bucket | Average |
@@ III-C Decision
Implementation details. We use the convolutional network-based diffusion policy [10]. We use DDIM [63] as the noise scheduler and use sample prediction instead of epsilon prediction for better high-dimensional action generation, with 100 timesteps at training and 10 timesteps at inference. We train 3000 epochs with batch size 128 for DP3 and all the baselines to ensure convergence for all the tasks.
@@ IV-A Experiment Setup
Expert demonstrations are collected with agents trained by reinforcement learning (RL) algorithms for all the domains except DexDeform, where we use human-teleoperated data. VRL3 [67] is used for Adroit; BAC [29] is used for MetaWorld; PPO [54] is used in all other domains. We generate successful trajectories with RL agents and ensure all imitation learning algorithms are using the same demonstrations. The success rates for experts are given in Appendix -C.
@@ IV-B Efficiency and Effectiveness
High accuracy. Summarized results are in Figure 1(a) and results for each domain are in Table I. We observe that DP3 achieves a success rate exceeding 90% in nearly 30 tasks, whereas Diffusion Policy does in fewer than 15 tasks. Additionally, DP3 did not record any task with a success rate below 10%, in contrast to Diffusion Policy, which had more than 10 tasks below 10%. Note that most of the tasks are only trained with 10 demonstrations.
@@
Learning efficiency. While we train all the algorithms for 3000 epochs to guarantee convergence, we observe that DP3 typically reaches convergence within approximately 500 epochs across all tasks, as illustrated in Figure 5. In contrast, Diffusion Policy tends to converge at a much slower pace or converge into sub-optimal results.
@@ IV-C Ablations
We select 6 tasks to conduct more ablation studies: Hammer (H), Door (D), Pen (P) from Adroit and Assembly (A), Basketball (B), Shelf Place (S) from MetaWorld. These tasks include both high-dimensional and low-dimensional control tasks, and each task only uses 10 demonstrations. We use the abbreviations of these tasks in the tables for simplicity.
Choice of 3D representations. In DP3, we deliberately select point clouds to represent the 3D scene. To compare different choices of 3D representations, we implement other 3D representations, including RGB-D, depth, and voxel. The RGB-D and depth images are processed using the same image encoder as Diffusion Policy, while voxel representations employ the VoxelCNN, as implemented in [7]. As demonstrated in Table IV, these alternative 3D representations fall short of DP3. Notably, both RGB-D and 2D images are outperformed by depth images, indicating that our tasks are more focused on low-level control rather than scene understanding.
| Repr. | H | D | P | A | B | S | Average |
@@
Gradually modifying a PointNet. To elucidate the performance disparity between DP3 Encoder and a commonly used point cloud encoder, e.g., PointNet, we gradually modify a PointNet to make it aligned with a DP3 Encoder. Through extensive experiments shown in Table VI, we identify that the T-Net and BatchNorm layers in PointNet are primary inhibitors to its efficiency. By omitting these two elements, PointNet attains an average success rate of , competitive to achieved by our DP3 Encoder. One plausible explanation for the T-Net is that our control tasks use the fixed camera and do not require feature transformations from the T-Net. We would explore the reason for failures of other encoders in the future.
@@
| Turnaround! | ✓ | ✗ | ✗ | ✓ | |
Design choices in DP3. Besides the 3D representations, the effectiveness of DP3 is contributed by several small design choices, as shown in Table VII. (a) Cropping point clouds helps largely improve accuracy; (b) Incorporating LayerNorm layers could help stabilize training across different tasks [23, 4]; (c) Sample prediction in the noise sampler brings faster convergence, also shown in Figure 7; (d) The projection head in DP3 Encoder accelerates the inference by projecting features to the lower dimension, without hurting accuracy; (e) Removing color channels ensures robust appearance generalization; (f) In low-dimensional control tasks, DPM-solver++ [34] as the noise sampler is competitive to DDIM, while DPM-solver++ could not handle high-dimensional control tasks well.
| Designs | H | D | P | A | B | S | Average |
@@
Training policies on demonstrations from different experts. We wanted to study whether the behavior modality of experts would affect the performance of imitation learning methods. To this end, we generate demonstrations with RL-trained agents and script policies respectively in MetaWorld tasks and train on these demonstrations. Summarized results are given in in Table VIII and detailed results are shown in Appendix -C. We could observe that a better expert could generally help train a better policy with imitation learning algorithms, while there still exists a gap between DP3 and the ground truth expert.
| MetaWorld 50 Tasks | Expert (Script) | Expert (BAC [29]) | |
| Ground Truth | |||
| Diffusion Policy | |||
| DP3 |
Different image resolutions for Diffusion Policy. In all experiments, we use the same resolution for DP3 and Diffusion Policy, while we still wanted to study whether improving the image resolution could help Diffusion Policy achieve better results. As shown in Table IX, we observe that a larger resolution such as and could improve Diffusion Policy in the Basketball task, while the success rates averaged across 6 tasks are still close. This again demonstrates the necessity of the 3D representations.
| Algorithm | H | D | P | A | B | S | Average |
| DP3 | |||||||
| Diffusion Policy (84) | |||||||
| Diffusion Policy (128) | |||||||
| Diffusion Policy (256) |
@@ V-B Effectiveness
Results for our real robot tasks are given in Table X. Consistent with our simulation findings, we observe in real-world experiments that DP3 could handle all tasks with high success rates, given only 40 demonstrations. Interestingly, we also observe that while both image-based and depth-based diffusion policies have comparatively low average accuracies, they exhibit distinct strengths in specific tasks. For instance, the image-based diffusion policy excels in the Drill task but fails entirely in Roll-Up. In contrast, the depth-based policy achieves a notable success rate of in Roll-Up.
@@ V-C Generalization
Spatial generalization. As illustrated in our motivating example, DP3 could better extrapolate in 3D space. We demonstrate this property in the real world, as shown in Table XI. We find that baselines fail to generalize to all test positions while DP3 succeed in 4 out of 5 trials.
![[Uncaptioned image]](x12.png)
@@
Appearance generalization. DP3 is designed to process point clouds without color information, inherently enabling it to generalize across various appearances effectively. As demonstrated in Table XII, DP3 consistently exhibits successful generalization to cubes of differing colors, while baseline methods could not achieve. It is noteworthy that the depth-based diffusion policy also does not incorporate color as input. However, due to its lower accuracy on the training object, the ability to generalize is also limited.
@@
Instance generalization. Achieving generalization across diverse instances, which vary in shape, size, and appearance, presents a significantly greater challenge compared to mere appearance generalization. In Table XIII, we demonstrate that DP3 effectively manages a wide range of everyday objects. This success can be primarily attributed to the inherent characteristics of point clouds. Specifically, the use of point clouds allows for policies that are less prone to confusion, particularly when these point clouds are downsampled. This feature significantly enhances the model’s ability to adapt to varied instances.
![[Uncaptioned image]](x14.png)
@@
View generalization. Generalizing image-based methods across different views is notably challenging [74], and acquiring training data from multiple views can be time-consuming and costly [81, 59]. We demonstrate in Table XIV that DP3 effectively addresses this generalization problem when the camera views are altered slightly. It is important to note that since the camera view is altered, we manually adjust the cropped space of the point clouds.
![[Uncaptioned image]](x15.png)
@@ V-D Safety Violation
In our real-world experiments, we surprisingly observe that image-based and depth-based diffusion policies often deliver unpredictable behaviors, which necessitates human termination to ensure robot safety. This situation is defined as safety violation. We compute the safety violation rate in our main real-world experiments, shown in Table XV. Interestingly, DP3 rarely violates the safety, showing that DP3 is a practical and hardware-friendly method for real robot learning.
![[Uncaptioned image]](x16.png)

### v2

@@ Abstract
Imitation learning provides an efficient way to teach robots dexterous skills; however, learning complex skills robustly and generalizablely usually consumes large amounts of human demonstrations. To tackle this challenging problem, we present 3D Diffusion Policy (DP3), a novel visual imitation learning approach that incorporates the power of 3D visual representations into diffusion policies, a class of conditional action generative models. The core design of DP3 is the utilization of a compact 3D visual representation, extracted from sparse point clouds with an efficient point encoder. In our experiments involving 72 simulation tasks, DP3 successfully handles most tasks with just 10 demonstrations and surpasses baselines with a 24.2% relative improvement. In 4 real robot tasks, DP3 demonstrates precise control with a high success rate of 85%, given only 40 demonstrations of each task, and shows excellent generalization abilities in diverse aspects, including space, viewpoint, appearance, and instance. Interestingly, in real robot experiments, DP3 rarely violates safety requirements, in contrast to baseline methods which frequently do, necessitating human intervention. Our extensive evaluation highlights the critical importance of 3D representations in real-world robot learning. Videos, code, and data are available on 3d-diffusion-policy.github.io.
@@ I Introduction
Safety. An interesting observation in our real-world experiments is that DP3 seldom gives erratic commands in real-world tasks, unlike baseline methods which often do and exhibit unexpected behaviors, posing potential damage to the robot hardware.
@@ II-A Diffusion Models in Robotics
Diffusion models, a category of generative models that progressively transform random noise into a data sample, have achieved great success in high-fidelity image generation [24, 64, 52, 63]. Owing to their impressive expressiveness, diffusion models have recently been applied in robotics, including in fields such as reinforcement learning [70, 2], imitation learning [10, 41, 51, 72, 65], reward learning [26, 39], grasping [71, 67, 62], and motion planning [53, 28]. In this work, we focus on representing visuomotor policies as conditional diffusion models, referred to as diffusion policies, following the framework established in [10, 41]. Unlike prior methods that primarily focus on images and states as conditions, we pioneer in incorporating 3D conditioning into diffusion policies.
@@ III-B Perception
| Algorithm Task | Hammer | Door | Pen | Assembly | Disassemble | Stick-Push | Laptop | Faucet | Toilet | Bucket | Average |
@@ III-C Decision
Implementation details. We use the convolutional network-based diffusion policy [10]. We use DDIM [63] as the noise scheduler and use sample prediction instead of epsilon prediction for better high-dimensional action generation, with 100 timesteps at training and 10 timesteps at inference. We train 1000 epochs for MetaWorld tasks and 3000 epochs for other simulated and real-world tasks, with batch size 128 for DP3 and all the baselines.
@@ IV-A Experiment Setup
Expert demonstrations. Human-teleoperated data is used in DexDeform; Script policies are used in MetaWorld; Trajectories for other domains are collected with agents trained by reinforcement learning (RL) algorithms, where we use VRL3 [68] is used for Adroit; BAC [29] is used for MetaWorld; PPO [54] is used in all other domains. We generate successful trajectories with RL agents and ensure all imitation learning algorithms are using the same demonstrations. The success rates for experts are given in Appendix -C.
@@ IV-B Efficiency and Effectiveness
High accuracy. Summarized results are in Figure 1(a) and results for each domain are in Table I. We observe that DP3 achieves a success rate 100% in nearly 20 tasks, whereas Diffusion Policy does in around 10 tasks. Additionally, we observe that Diffusion Policy records more than 20 tasks with a success rate lower than 30% given a small set of demonstrations, while DP3 could generally achieve higher accuracy.
@@
Learning efficiency. While we train enough epochs to guarantee convergence, we observe that DP3 typically reaches convergence within approximately 500 epochs across all tasks, as illustrated in Figure 5. In contrast, Diffusion Policy tends to converge at a much slower pace or converge into sub-optimal results.
@@ IV-C Ablations
We select 6 tasks to conduct more ablation studies: Hammer (H), Door (D), Pen (P) from Adroit and Assembly (A), Disassemble (DA), Stick-Push (SP) from MetaWorld. These tasks include both high-dimensional and low-dimensional control tasks, and each task only uses 10 demonstrations. We use the abbreviations of these tasks in the tables for simplicity.
Choice of 3D representations. In DP3, we deliberately select point clouds to represent the 3D scene. To compare different choices of 3D representations, we implement other 3D representations, including RGB-D, depth, and voxel. The RGB-D and depth images are processed using the same image encoder as Diffusion Policy, while voxel representations employ the VoxelCNN, as implemented in [7]. As demonstrated in Table IV, these alternative 3D representations fall short of DP3. Notably, RGB-D and depth images are very close and not comparable to point clouds, indicating that the proper usage of depth information is essential.
| Repr. | H | D | P | A | DA | SP | Average |
@@
Gradually modifying a PointNet. To elucidate the performance disparity between DP3 Encoder and a commonly used point cloud encoder, e.g., PointNet, we gradually modify a PointNet to make it aligned with a DP3 Encoder. Through extensive experiments shown in Table VI, we identify that the T-Net and BatchNorm layers in PointNet are primary inhibitors to its efficiency. By omitting these two elements, PointNet attains an average success rate of , competitive to achieved by our DP3 Encoder.One plausible explanation for the T-Net is that our control tasks use the fixed camera and do not require feature transformations from the T-Net. Further replacing high-dimensional features with a lower-dimensional one would not hurt the performance much () but increase the speed. We would explore the reason for the failures of other encoders in the future.
@@
| Turnaroud! | ✓ | ✗ | ✗ | ✓ | |
Design choices in DP3. Besides the 3D representations, the effectiveness of DP3 is contributed by several small design choices, as shown in Table VII. (a) Cropping point clouds helps largely improve accuracy; (b) Incorporating LayerNorm layers could help stabilize training across different tasks [23, 4]; (c) Sample prediction in the noise sampler brings faster convergence, also shown in Figure 7; (d) The projection head in DP3 Encoder accelerates the inference by projecting features to the lower dimension, without hurting accuracy; (e) Removing color channels ensures robust appearance generalization; (f) In low-dimensional control tasks, DPM-solver++ [34] as the noise sampler is competitive to DDIM, while DPM-solver++ could not handle high-dimensional control tasks well empirically.
| Designs | H | D | P | A | DA | SP | Average |
@@ V-B Effectiveness
Results for our real robot tasks are given in Table VIII. Consistent with our simulation findings, we observe in real-world experiments that DP3 could handle all tasks with high success rates, given only 40 demonstrations. Interestingly, we also observe that while both image-based and depth-based diffusion policies have comparatively low average accuracies, they exhibit distinct strengths in specific tasks. For instance, the image-based diffusion policy excels in the Drill task but fails entirely in Roll-Up. In contrast, the depth-based policy achieves a notable success rate of in Roll-Up.
@@ V-C Generalization
Spatial generalization. As illustrated in our motivating example, DP3 could better extrapolate in 3D space. We demonstrate this property in the real world, as shown in Table IX. We find that baselines fail to generalize to all test positions while DP3 succeed in 4 out of 5 trials.
@@
Appearance generalization. DP3 is designed to process point clouds without color information, inherently enabling it to generalize across various appearances effectively. As demonstrated in Table X, DP3 consistently exhibits successful generalization to cubes of differing colors, while baseline methods could not achieve. It is noteworthy that the depth-based diffusion policy also does not incorporate color as input. However, due to its lower accuracy on the training object, the ability to generalize is also limited.
@@
Instance generalization. Achieving generalization across diverse instances, which vary in shape, size, and appearance, presents a significantly greater challenge compared to mere appearance generalization. In Table XI, we demonstrate that DP3 effectively manages a wide range of everyday objects. This success can be primarily attributed to the inherent characteristics of point clouds. Specifically, the use of point clouds allows for policies that are less prone to confusion, particularly when these point clouds are downsampled. This feature significantly enhances the model’s ability to adapt to varied instances.
@@
View generalization. Generalizing image-based methods across different views is notably challenging [75], and acquiring training data from multiple views can be time-consuming and costly [82, 59]. We demonstrate in Table XII that DP3 effectively addresses this generalization problem when the camera views are altered slightly. It is important to note that since the camera view is altered, we manually adjust the cropped space of the point clouds.
@@ V-D Safety Violation
In our real-world experiments, we surprisingly observe that image-based and depth-based diffusion policies often deliver unpredictable behaviors, which necessitates human termination to ensure robot safety. This situation is defined as safety violation. We compute the safety violation rate in our main real-world experiments, shown in Table XIII. Interestingly, DP3 rarely violates the safety, showing that DP3 is a practical and hardware-friendly method for real robot learning.